A certain particle falls under gravity in a resisting medium whose resistance varies with velocity.Find the relation between distance and velocity if the initially the particle starts from rest.


1
Expert's answer
2021-08-04T16:34:40-0400

"\textsf{If } v \textsf{ is the velocity of the particle falling under}\\
\textsf{gravity, then the equation of motion of the moving}\\
 \textsf{particle by Newton's second law is given by}\\ \hspace{0.1cm}\\
mg-mR=ma\\ \hspace{0.1cm}\\
\textsf{where } R\textsf{ represents the resistance of the }\\ 
\textsf{resisting medium}\\
\hspace{0.1cm}\\
\textsf{ Since}\textsf{ R }\textsf{is proportional to } v,\\\hspace{0.1cm}\\
R=kv\\\hspace{0.1cm}\\
mg-mkv=ma\\
g-kv=a\\
\hspace{0.1cm}\\
a=\frac{\mathrm{d}v}{\mathrm{d}t}\\
\textsf{Using chain rule,}\\
a=\frac{\mathrm{d}v}{\mathrm{d}s}\times\frac{\mathrm{d}s}{\mathrm{d}t}\\
a=\frac{\mathrm{d}v}{\mathrm{d}s}\times v\\
a=v\frac{\mathrm{d}v}{\mathrm{d}s}\\
\hspace{0.1cm}\\
\textsf{Hence,}\\
g-kv=v\frac{\mathrm{d}v}{\mathrm{d}s}\\
\mathrm{d}s=(\frac{v}{g-kv})\mathrm{d}v\\
\textsf{By separating into partial fractions, we have}\\
\mathrm{d}s=(\frac{-1}k+\frac{g}{k(g-kv)})\mathrm{d}v\\
\mathrm{d}s=\frac{-\mathrm{d}v}{k}+\frac{g}{k}\frac{\mathrm{d}v}{g-kv}\\\hspace{0.1cm}\\
\textsf{Integrating, we have,}\\
\hspace{0.1cm}\\
s+c=\frac{-v}{k}+\frac{g}{k}(\frac{-1}{k})ln(g-kv)\\
s+c=\frac{-v}{k}-\frac{g}{k^2}ln(g-kv)\\
\hspace{0.1cm}\\
\textsf{Initially, } s=0, v=0.\, \textsf{Hence,}\\\hspace{0.1cm}\\
c=\frac{-g}{k^2}ln(g)\\\hspace{0.1cm}\\
\textsf{Hence,}\hspace{0.1cm}\\
s-\frac{g}{k^2}ln(g)=\frac{-v}{k}-\frac{g}{k^2}ln(g-kv)\\
s=-\frac{v}{k}-\frac{g}{k^2}ln(\frac{g-kv}{g})\\
s=-(\frac{v}{k}+\frac{g}{k^2}ln(\frac{g-kv}{g})"